Job Costs Using Activity-Based Costing

Heitger Company is a job-order costing firm that uses activity-based costing to apply overhead to jobs. Heitger identified three overhead activities and related drivers. Budgeted information for the year is as follows:

Activity Cost Driver Amount of Driver
Materials handling $71,100 Number of moves 3,000
Engineering 105,300 Number of change orders 6,500
Other overhead 238,500 Direct labor hours 45,000

Heitger worked on four jobs in July. Data are as follows:

Job 13-43 Job 13-44 Job 13-45 Job 13-46
Beginning balance $20,800 $17,200 $3,000 $0
Direct materials $6,900 $8,600 $11,400 $9,800
Direct labor cost $850 $1,000 $1,640 $110
Job 13-43 Job 13-44 Job 13-45 Job 13-46
Number of moves 40 46 28 8
Number of change orders 31 44 18 18
Direct labor hours 850 1,000 1,640 110

By July 31, Jobs 13-43 and 13-44 were completed and sold. Jobs 13-45 and 13-46 were still in process.

Required:

1. Calculate the activity rates for each of the three overhead activities. Round all activity rates to the nearest cent.

Materials handling rate $ per move
Engineering rate $ per change order
Other overhead rate $ per direct labor hour

2. Prepare job-order cost sheets for each job showing all costs through July 31.

When required, round your answers to the nearest dollar. If an amount is zero, enter "0".

Heitger Company
Job-Order Cost Sheets
Job 13-43 Job 13-44 Job 13-45 Job 13-46
Balance, July 1 $ $ $ $
Direct materials
Direct labor cost
Materials handling
Engineering
Other overhead
Total $ $ $ $

3. Calculate the balance in Work in Process on July 31.
$

4. Calculate the cost of goods sold for July.
$

5. What if Job 13-46 required no engineering change orders? What is the difference in the new cost of Job 13-46?

$    

How would the cost of the other jobs be affected?
Answer #1
1 Activity rate=Cost of activity/Amount of driver
Activity Cost Driver Amount of driver Activity rate
a b a/b
Materials handling 71100 Number of moves 3000 23.7 Per move
Engineering 105300 Number of change orders 6500 16.2 Per change order
Other overhead 238500 Direct labor hours 45000 5.3 Per direct labor hour
2 Job-Order cost sheets
Job 13-43 Job 13-44 Job 13-45 Job 13-46
Balance, July 1 20800 17200 3000 0
Direct materials 6900 8600 11400 9800
Direct labor cost 850 1000 1640 110
Materials handling 948 1090 664 190
(Number of moves*Activity rate) (40*23.7) (46*23.7) (28*23.7) (8*23.7)
Engineering 502 713 292 292
(Number of change orders*Activity rate) (31*16.2) (44*16.2) (18*16.2) (18*16.2)
Other overhead 4505 5300 8692 583
(Direct labor hours*Activity rate) (850*5.3) (1000*5.3) (1640*5.3) (110*5.3)
Total 34505 33903 25687 10974
3 Balance in work in process on July 31=Cost of jobs still in process=Cost of Job 13-45 and 13-46=25687+10974=$ 36661
4 Cost of goods sold=Cost of jobs sold=Cost of job 13-43 and 13-44=34505+33903=$ 68408
5 New cost:
Job 13-46
Balance, July 1 0
Direct materials 9800
Direct labor cost 110
Materials handling 190
(Number of moves*Activity rate) (8*23.7)
Engineering 0
(Number of change orders*Activity rate) (0*16.2)
Other overhead 583
(Direct labor hours*Activity rate) (110*5.3)
Total 10683
Cost of other jobs will not be affected by this